- 2/28-3/5: Fixing PRMs and SRMs
- 3/2-5: Re-alignment
- 3/6-7: Closing
1/25-2/20: Yend Pcal finalization
- 2/25-3/4: Yend WAB finalization
- 3/5-6: Closing
- 3/6~ : Yend pumpdown 3/6~
- 3/7,8: leak test
2/11-2/14: Xend NAB installation
2/21: Removing FC, Payload release
2/27: ETMX release
- 2/23-3/10: Xend Pcal finalization 1
- Periscope installation
- Beam position by IR and Green
- 3/5: ETMX aligned
- 3/5-3/10: Xend Pcal finalization 2
- 3/11,12 Xend WAB finalization
- 3/19,20, Heatlink etc.
- 3/22 Pumping down
- Leak test
- 3/11~ : Night work starts
- 2/18-3/22 DRMI commissioning
- 3/11-4/5 Y arm commissioning
- 3/25-4/5 FPMI commissioning
Report of work
- PR3 was tuned once but oplev was gone again. Needs tuning one more day.
- Type A: ETMX seems not healthy. To be fixed the next week.
- Type B:Offset on IP of BS was seen but not real motion.
- IOO: Beam damp for misalignment of PRM seems not working well. Mirror cannot be tilted well.
- CAL: Pcal alignment was done. AOS will finish finalization. We need to confirm scheduling at sitework meeting or commissioning meeting.